# A Novel Mathematical Framework for Objective Characterization of Ideas through Vector Embeddings in LLM

### v2

@@ 2. Idea Generation using CAI
0.2 2.2. Generating Ideas Using AOC
The problem statements serve as prompts for the CAI system, guiding the generation of relevant ideas. For this study, we used the following six distinct problem statements.
@@ 2 Queue Companion Stool Provide temporary seating Stool Allows rest without losing place in the queue
0.3 2.3. Advantages of Using CAI for Idea Generation
@@
3. Characteristics of Idea Exploration
@@
0.4 3.1. Concept of Idea Space
@@
0.4.1 3.1.1. Cognitive Processes Involved in Exploration of Idea Space
@@
0.4.2 3.1.2. Challenges in the Exploration of Idea Space
@@
0.5 3.2. Measures for Quantifying Exploration of Idea Space
@@
0.6 3.3. Classification of Idea Space
@@
4. Embeddings for Idea Representation
@@
0.9 4.3. Properties and Types of Embeddings
@@
0.10 4.4. Types of Embedding Models
@@
0.10.1 4.4.1. Static Embedding Models
@@
0.10.2 4.4.2. Dynamic Embedding Models
@@
0.11 4.5. Process of Conversion of Text to Embedding Vectors
@@
0.13 4.7. Demonstrating the Validity of Embeddings
@@
5. Requirements for Idea Embedding
@@
1 6. Empirical Exploration of Idea Embedding
@@
1.2 6.2. Meaningfulness of Idea Embeddings (RQ1.1)
@@
1.3 6.3. Usefulness of Idea Embeddings (RQ1.2)
@@
1.4 6.4. Distribution of Ideas in the Idea Space (RQ2.1)
@@
An effective ideation exercise would have a set of ideas that cover a large region in the idea space. This means that the mean distance between two ideas in that space is large. We refer to this as a Sparse distribution. When the ideas are all diverse, no two ideas would belong to a cluster because each cluster represents a set of similar ideas. However, in the presence of clustering, the mean distance between the ideas in different clusters is usually different. Also, the larger the number of clusters, the greater the number of distinct ideas. The distinctness of ideas within each such cluster can be categorized based on their distribution into Sparse (Small No. of ideas over Larger Area), Moderate (Large/Small No. of ideas over Larger/Smaller Area respectively) and Dense (Large No. of ideas over Smaller Area). A sparse cluster is considered locally more effective than moderate, and similarly, a moderate is more effective than dense. We define Idea Sparsity (IS) to capture these distribution as follows:
@@
where $N_{i}$ is the Number of Ideas in a cluster and $A_{c}$ is the Area of the cluster computed as the area of the convex hull of the points in that cluster. If the idea sparsity is similar across all the clusters, we say that the idea exploration is uniform. In that case, in a spider plot (Figure 10), where each spoke corresponds to a cluster and the length of the spoke corresponds to idea sparsity, a uniform exploration would appear as a regular convex polygon. The larger the polygon, the more number of ideas are distributed uniformly over a larger area within each cluster.
Each cluster in an idea space is distributed over an area. Ideally, we desire to have clusters that are sparsely distributed from each other since they represent more distinct sets of ideas compared to clusters that are close to each other, as evidenced by the expert opinion in the section –. In order to quantify the nature of the distribution of clusters in the idea space, we define Cluster Sparsity (CS) as follows:
@@
where $N_{c}$ is the number of clusters in an idea space, $A_{i}$ is the area of individual clusters, and $A_{t}$ is the total area of the idea space computed as the area of the convex hull of all the points in that idea space. An exploration where the clusters are well separated would have high cluster sparsity (desirable), and a compact set of clusters will give low cluster sparsity (undesirable).
DBSCAN clusters potentially produce a default cluster with points that could not be included in any of the well-defined clusters; this is usually referred to as the noise cluster. This is the one which is not indicated with a bounding curve. The points in this cluster are important as the associated ideas are distinct from clusters of similar ideas. We compute Idea Sparsity for all clusters, including the noise cluster. However, we observe that (a) noise clusters do not significantly extend the region of explored ideas, and (b) the convex bounding curve for the noise cluster inevitably overlaps multiple identified clusters. Hence, for the computation of cluster sparsity, we discount the noise cluster.
In terms of cluster and idea sparsities, an effective idea exploration would have a high cluster sparsity with a regular polygon of large size for idea sparsity. In Figure 10, the brightness of the bounded polygon of idea sparsity represents the cluster sparsity.
@@
1.5 6.5. Dispersion of Ideas in the Idea Space (RQ2.2)
Here, we assess whether ideas spanned all dimensions of the idea space effectively. Using the measures described above, comprehensive coverage of all the dimensions of the embedding cannot be assessed as it is based on a 2D map of the original embeddings. To assess the dispersion of ideas, which refers to the extent to which the ideas are spanned in the higher-dimensional space, Principal Component Analysis (PCA) was adopted. PCA provides eigenvectors that describe the direction and relative magnitude of dominant explorations. An effective exploration of all dimensions of idea space would give comparable eigenvalues of reasonably large magnitude. For the assessment of the comprehensiveness of different ideation exercises, the eigenvalues were computed using PCA for each idea set and sorted in descending order. The results are presented in Figure 11(a) and Figure 11(b). It was observed that for all the dimensions, the idea space is explored more or less uniformly except for the first few dominant eigen directions. The maximum eigenvalue observed was 300, although all eigenvalues eventually saturated in the range of 30 to 50. Hence, the comparison of exploration below uses the first 10 eigenvalues only.
@@
7. Discussion

### v3

@@ 2.3 Idea Representation
2.4 Idea Generation
The user interface of the tool was designed and developed by the authors as shown in Figure 2 with labels for different fields. The DesignerAI is now integrated into a standalone software application that has not yet been publicly released. The preliminary testing from our earlier study (Sankar2024) has shown promising results in its ability to assist designers in generating a variety of ideas. In the same study, the authors developed a structure for the input prompts and output responses. The creativity of the ideas generated by our tool is controlled with a field marked as ’creativity slider’, which adjusts the parameter called temperature; a higher temperature setting produces more diverse outputs, while a lower temperature setting produces focused and predictable ideas.
The problem statements serve as input prompts for the CAI system, guiding the generation of relevant ideas. It was established in Sankar2024 that the statements of the machine-generated ideas were valid and human-like to the experts and faired significantly better vis-a-vis human-generated ideas in terms of novelty, fluency and variety. That study used the following six problem statements derived from common day-to-day experiences.
@@ 2 Queue Companion Stool Provide temporary seating Stool Allows rest without losing place in the queue
2.5 Advantages of Using CAI for Idea Generation
@@
Characteristics of Idea Exploration
@@
2.6 Concept of Idea Space
@@
2.6.1 Cognitive Processes Involved in Exploration of Idea Space
@@
2.6.2 Challenges in the Exploration of Idea Space
@@
2.7 Measures for Quantifying Exploration of Idea Space
@@ 2.7.2 Distribution
2.8 Classification of Idea Space
@@
Embeddings for Idea Representation
@@ 2.10 Mechanism of Embeddings
2.11 Properties and Types of Embeddings
@@
2.12 Types of Embedding Models
@@
2.12.1 Static Embedding Models
@@
2.12.2 Dynamic Embedding Models
@@
2.13 Process of Conversion of Text to Embedding Vectors
@@ 2.14 Contribution to CAI Behavior
2.15 Demonstrating the Validity of Embeddings
@@ 2.15.2 Sentence Embedding
Requirements of Idea Embedding for Designers
@@
3 Empirical Exploration of Idea Embedding
@@ 3.1 Methodology
3.2 Meaningfulness of Idea Embeddings (RQ1.1)
@@ 3.2.1 Results and Inference
3.3 Usefulness of Idea Embeddings (RQ1.2)
@@ 3.3.1 Results and Inference
3.4 Distribution of Ideas in the Idea Space (RQ2.1)
@@
Figure 11: Representative diagram illustrating four types of idea clusters characterized by varying spatial distributions—Sparse, Moderate (Large Spread), Moderate (Small Spread), and Dense. Each cluster is visualized with distinct colors and symbols, enclosed by a dotted convex hull boundary. The corresponding Idea Sparsity (IS) values quantify the spatial spread of the ideas within each cluster.
(a) Tight-Knit Clustering Pattern
(b) Loosely-Woven Clustering Pattern
Figure 12: Representative diagram illustrating two clustering scenarios—Tight-Knit Clustering and Loosely-Woven Clustering—with identical intra-cluster spread across five distinct, non-overlapping clusters. Dotted convex hulls outline each cluster, while the Cluster Sparsity (CS) metric captures the degree of global separation. A higher CS in the loosely-woven configuration signifies more distinct and spatially dispersed idea groups. Note: The corresponding clusters are of the same size on both plots.
3.4.1 Idea Sparsity: Quantifying Local Distribution of Ideas within a Cluster
An effective ideation exercise would have a set of ideas that cover a large region in the idea space. This means that the mean distance between two ideas in that space is large. We refer to this as a Sparse distribution. When the ideas are all diverse, no two ideas would belong to a cluster because each cluster represents a set of similar ideas. However, in the presence of clustering, the mean distance between the ideas in different clusters is usually different. Also, the larger the number of clusters, the greater the number of distinct ideas. The distinctness of ideas within each such cluster can be categorized based on their distribution into Sparse (Small No. of ideas over Larger Area), Moderate (Large/Small No. of ideas over Larger/Smaller Area respectively) and Dense (Large No. of ideas over Smaller Area) (as shown in Figure 11). A sparse cluster is considered locally more effective than moderate, and similarly, a moderate cluster is more effective than dense. We define Idea Sparsity (IS) as a measure of how sparsely the ideas are distributed within a cluster in the idea space. It captures the density-independent spatial spread of ideas, which reflects the degree of local diversity among semantically similar ideas.
The equation used to calculate the Idea Sparsity (IS) for a given cluster i is as follows:
@@ 3.4.1 Idea Sparsity: Quantifying Local Distribution of Ideas within a Cluster
where,
$N_{i}$ is the Number of Ideas in a cluster
$A_{c}$ is the Area of the cluster computed as the area of the convex hull of the points in that cluster.
If the idea sparsity is similar across all the clusters, we say that the idea exploration is uniform. In that case, in a spider plot (Figure 13), where each spoke corresponds to a cluster and the length of the spoke corresponds to idea sparsity, a uniform exploration would appear as a regular convex polygon. The larger the polygon, the more number of ideas are distributed uniformly over a larger area within each cluster.
3.4.2 Cluster Sparsity (CS): Quantifying Global Separation between Clusters in the Idea Space
Each cluster in an idea space is distributed over an area. Ideally, we desire to have clusters that are sparsely distributed from each other (Loosely-woven clustering - Figure 12(b)) since they represent more distinct sets of ideas compared to clusters that are close to each other (Tight-knit clustering - Figure 12(a)), as evidenced by the expert opinion in the section 3.2. In order to quantify the nature of the distribution of clusters in the idea space, we define Cluster Sparsity (CS) as a measure of how well-separated the clusters are in the overall idea space. It reflects the global distinctness between groups of ideas, each representing different semantic themes.
The Cluster Sparsity (CS) is calculated using the equation as follows:
@@ 3.4.2 Cluster Sparsity (CS): Quantifying Global Separation between Clusters in the Idea Space
where,
$N_{c}$ is the number of clusters in an idea space,
$A_{i}$ is the area of individual clusters,
$A_{t}$ is the total area of the idea space computed as the area of the convex hull of all the points in that idea space.
An exploration where the clusters are well separated would have high cluster sparsity (desirable), and a compact set of clusters would give low cluster sparsity (undesirable). In Figure 13, the brightness of the bounded polygon represents the cluster sparsity.
DBSCAN clusters potentially produce a default cluster with points that could not be included in any of the well-defined clusters; this is usually referred to as the noise cluster. This is the one that is not indicated with a bounding curve. The points in this cluster are important as the associated ideas are distinct from clusters of similar ideas. We compute Idea Sparsity for all clusters, including the noise cluster. However, we observe that (a) noise clusters do not significantly extend the region of explored ideas, and (b) the convex bounding curve for the noise cluster inevitably overlaps multiple identified clusters. Hence, for the computation of cluster sparsity, we discount the noise cluster.
3.4.3 Results and Inference
In terms of cluster and idea sparsities, an effective idea exploration would have a high cluster sparsity with a regular polygon of large size for idea sparsity.
@@ 3.4.3 Results and Inference
3.5 Dispersion of Ideas in the Idea Space (RQ2.2)
Here, we assess whether ideas spanned all dimensions of the idea space effectively. Using the measures described above, comprehensive coverage of all the dimensions of the embedding cannot be assessed as it is based on a 2D map of the original embeddings. To assess the dispersion of ideas, which refers to the extent to which the ideas are spanned in the higher-dimensional space, Principal Component Analysis (PCA) was adopted. PCA provides eigenvectors that describe the direction and relative magnitude of dominant explorations. An effective exploration of all dimensions of idea space would give comparable eigenvalues of reasonably large magnitude. For the assessment of the comprehensiveness of different ideation exercises, the eigenvalues were computed using PCA for each idea set and sorted in descending order.
3.5.1 Results and Inferences
The results are presented in Figure 14(a) and Figure 14(b). It was observed that for all the dimensions, the idea space is explored more or less uniformly except for the first few dominant eigen directions. The maximum eigenvalue observed was 300, although all eigenvalues eventually saturated in the range of 30 to 50. Hence, the comparison of exploration below uses the first 10 eigenvalues only.
@@ 3.5.1 Results and Inferences
4 Evaluation through existing AI-Based Methods
Automated evaluation techniques leveraging Artificial Intelligence (AI) have recently emerged as prominent tools for assessing ideas. To explore their applicability in supporting novice designers during ideation, we employed five AI-based methods—namely, GPT-based evaluation, OCSAI (Open Creativity Scoring with Artificial Intelligence), AIDE (AI for Design Evaluation), CLAUS (Cross-Lingual Alternate Uses Scoring), and SemDis (Semantic Distance). These methods predominantly evaluate individual ideas based on a predefined scale (1 to 5), assessing factors such as originality or novelty relative to previously existing ideas. To assess their suitability, the same set of 100 previously generated ideas for each of six distinct product design problems (PS1 through PS6) was independently applied to each of five methods.
4.1 Results and Observations
The results of these evaluations were represented through two primary visualization techniques—box and whisker plots (depicting mean, median, standard deviation, Q1, and Q3) and bar plots with individual ratings, mean scores, and standard deviation markers. The plots are given in the appendix. Across the six problem statements, mean scores varied marginally within a relatively narrow range for each method, indicating a moderate degree of consistency at the aggregated statistical level. For instance, the AIDE method yielded a mean novelty score of 3.83 with standard deviations (SD) of 0.31, whereas CLAUS ratings ranged slightly lower (mean score of 3.43, SD of 0.2). Similarly, GPT and SemDis exhibited comparable average novelty scores, maintaining mean values of 3.32 and 3.09. However, OCSAI scores were notably lower, with a mean rating of 2.28 and higher variability (SD of 0.7). Despite similar statistical summaries (mean and SD), significant differences in score distributions and individual assessments were apparent. GPT-based assessments, for instance, displayed wider variability with frequent extreme values, whereas CLAUS and AIDE exhibited narrow distribution clusters. SemDis and OCSAI distributions displayed varied spread with moderate consistency.
Figure 15: Bar plot comparing the novelty ratings of ideas for Problem Statement 1 as assessed by five different AI-based methods
4.2 Limitations of the existing AI-based assessment Methods
Despite the systematic approach and the quantitative nature of these methods, these AI-based methods primarily provide absolute ratings of individual idea novelty, but do not offer mechanisms for comparative or relational evaluation among ideas, nor provide any insights into the nature of the ideation process that led to the generation of these ideas. Importantly, such approaches provide limited insights into the relative diversity or distinctiveness of ideas within a given set. For instance, two ideas scored equally high for novelty by AIDE or GPT might be conceptually very similar, providing little added value to designers looking for distinctly different ideas. Moreover, the assessment outcomes were inconsistent across different methods. Ideas receiving high ratings from GPT-based assessment frequently differed from those highly rated by OCSAI, CLAUS, or SemDis across all ideas. This inconsistency reflects methodological variations in criteria, model training, and sensitivity to input nuances, complicating the designers’ task of meaningfully interpreting these results to confidently shortlist promising ideas. Furthermore, none of these methods produced ratings below a certain threshold (e.g., below 2), reducing their discriminative power and limiting their utility for designers seeking genuinely innovative and diverse idea pools. These methods, while effectively assigning novelty ratings of ideas, fail to adequately characterize the ideation process and differentiate ideas based on conceptual diversity.
4.3 Advantages of our Proposed Embedding-based Clustering Method
Recognizing the shortcomings inherent in existing AI-based evaluation techniques, our proposed embedding-based clustering method addresses these critical gaps. Rather than assigning individual, isolated scores, our method leverages high-dimensional embeddings to represent ideas spatially. Through clustering techniques, it explicitly focuses on identifying distinctive groups of ideas, thus highlighting conceptual diversity and uniqueness within the idea pool.
4.3.1 Comparative Evaluation for a Large Set of Ideas
In contrast to existing AI-based methods that evaluate ideas in isolation, the embedding-based approach evaluates ideas collectively in relation to each other. Ideas are represented as points in a high-dimensional vector space, and semantic similarity or distance among ideas is computed systematically. This representation enables designers to visualize the relative positioning of ideas, making it significantly easier to recognize conceptual redundancy and diversity within the idea set. As a result, designers can make informed and strategic choices, selecting ideas not only for their individual novelty but also for their comparative distinctiveness and collective representativeness.
4.3.2 Enhanced Support for Novice Designers
One main benefit of our embedding-based approach is its suitability in assisting novice designers who typically lack expertise in effectively managing and analyzing large ideation sets. Unlike existing methods, which leave designers to interpret and select ideas based solely on individual novelty scores, embedding-based clustering provides an intuitive, visual, and structured representation of ideas. By mapping semantic relationships clearly, it equips novice designers with the capacity to quickly identify distinct clusters of ideas, thereby significantly reducing cognitive load and decision-making complexity associated with selection tasks. Moreover, clustering inherently aids in the identification of ”hidden gems”, the ideas that might otherwise be overlooked due to less compelling isolated novelty ratings but that distinctly stand out within specific clusters, enriching diversity within the selected shortlist. Such strategic selection from diverse clusters ensures comprehensive exploration of the ideation space, enhancing overall design innovation potential.
4.3.3 Transparent, Scalable, and Reproducible Outcomes
Another advantage of our method lies in its transparency and reproducibility. AI-based scoring methods often lack clarity in their reasoning, making it challenging to justify scores or replicate evaluations consistently across contexts. In contrast, embedding-based clustering provides transparent reasoning grounded in semantic relationships and geometric structures, which remain stable and reproducible across multiple assessments. The resultant clusters can be quantitatively and qualitatively justified, lending greater validity and reliability to the assessment process. Furthermore, embedding-based methods scale efficiently to handle large volumes of ideas without sacrificing accuracy or interpretability, thus fitting well into realistic industrial design scenarios characterized by extensive ideation outputs.
In summary, while existing methods effectively quantify individual idea novelty, their isolated and non-comparative evaluations severely limit their practical utility for novice designers tasked with shortlisting promising and diverse ideas from large pools. Our embedding-based clustering approach fundamentally addresses these issues by providing comparative, interpretable, and visually intuitive evaluations of idea sets. The method empowers novice designers by highlighting conceptual uniqueness and relational diversity, thereby supporting more informed, strategic, and efficient idea selection decisions. Consequently, our approach significantly outperforms existing AI-based methods in providing practical, scalable, and meaningful assistance during the crucial selection phase of the design process.
Discussion
@@ Conclusion
Appendix A Evaluation Results: Novelty Rating Distributions of AI-Based Assessment Methods
A.1 Bar Plots
The subsequent bar plot offers a visual representation of each novelty rating within the output of each AI assessment method. This allows for a more granular understanding of the rating patterns.
(a) GPT-based Assessment
(b) OCSAI-based Assessment
(c) AIDE-based Assessment
(d) CLAUS-based Assessment
(e) SemDis-based Assessment
Figure 17: Bar plots of novelty ratings for 100 ideas across 6 problem statements
A.2 Box and Whisker Plots
The box and whisker plots shown below provide a comparative summary of the distribution of novelty scores assigned by the different AI-based assessment methods. Key statistical features such as the mean, standard deviation, median, quartiles, and range are clearly depicted.
(a) GPT-based Assessment
(b) OCSAI-based Assessment
(c) AIDE-based Assessment
(d) CLAUS-based Assessment
(e) SemDis-based Assessment
Figure 18: Box and whisker plots illustrating the novelty ratings generated by five distinct AI-based assessment methods.